40m QIL Cryo_Lab CTN SUS_Lab CAML OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Oct 26 19:10:59 2015, gautam, Update, Green Locking, AUX PDH loop characterization 
    Reply  Tue Dec 8 23:24:08 2015, gautam, Update, Green Locking, Y end laser (Lightwave) PZT calibration Ycalib_8Dec.pdf
       Reply  Sun Dec 13 21:55:28 2015, gautam, Update, Green Locking, Y end laser (Lightwave) PZT calibration IMG_5972.JPGServoY_TF_13Dec2015.pdfDatanCode.zipPSD_916Hz.pdf
          Reply  Mon Dec 14 16:27:11 2015, gautam, Update, Green Locking, Y-end AUX PDH noise breakdown YAUX_NB_Dec2015.pdfPDH_errSig_Calib.pdf
             Reply  Mon Jan 4 16:45:11 2016, gautam, Update, Green Locking, Y-end AUX PDH noise breakdown ErrSigBreakdown.pdfcontrolSigBreakdown.pdfYEnd_PDH_OLTF.pdf
          Reply  Mon Jan 4 16:09:54 2016, gautam, Update, Green Locking, Y end laser (Lightwave) PZT calibration choosingExcFreqs.pdflaserPZTcalib.pdf
Message ID: 11907     Entry time: Mon Jan 4 16:45:11 2016     In reply to: 11879
Author: gautam 
Type: Update 
Category: Green Locking 
Subject: Y-end AUX PDH noise breakdown 

Summary:

I've re-measured the noise breakdown for the Y-end AUX PDH system. Spectra are attached. I've also measured the OLTF of the PDH loop, from which the UGF appears to be ~8.5kHz. 

Discussion:

As Eric and Koji pointed out, the spectra uploaded here were clearly wrong as there were breaks in the spectra between decades of frequency. I redid the measurements, this time being extra careful about impedance mismatch effects. All measurements were made from the monitor points on the PDH box, which according to the schematic found here, have an output impedance of 49.9 ohms. So for all measurements made using the SR785 which has an input impedance of 1Mohm, or those which had an SR560 in the measurement chain (also high input impedance), I terminated the input with a 50ohm terminator so as to be able to directly match up spectra measured using the two different analyzers. I'm also using my more recent measurement of the actuator gain of the AUX laser to convert the control signal from V/rtHz to Hz/rtHz in the plotted spectra. 

As a further check, I locked the IR to the Y-arm by actuating on MC2, and took the spectrum of the Y-arm mirror motion using the C1CAL model. We expect this to match up well with the in-loop control signal at low frequencies. However, though the shapes seem consistent in Attachment #2 (light orange and brown curves), I seem to be off by a factor of 5- not sure why. In converting the Y-arm mirror motion spectrum from m/rtHz to Hz/rtHz, I multiplied the measured spectrum by \frac{3.907*10^6}{0.5*532*10^{-9}}, which I think is the correct conversion factor (FSR/(0.5*wavelength))?

Attachment 1: ErrSigBreakdown.pdf  156 kB  | Show | Hide all | Show all
Attachment 2: controlSigBreakdown.pdf  190 kB  | Show | Hide all | Show all
Attachment 3: YEnd_PDH_OLTF.pdf  99 kB  | Hide | Hide all | Show all
YEnd_PDH_OLTF.pdf
ELOG V3.1.3-